[figure: 9.7 cm]

MAGNETIC STRUCTURES IN ULTRATHIN FILMS: SIMULATION

J.-C.S. Lévy^{a}, A. Ghazali^{b} and E.Yu. Vedmedenko^{c} aLPTMC, case 7020, Université Paris 7, 75251 Paris 5, France bGPS, CNRS, Université Paris 7 & Paris 6, 75251 Paris 5, France cMPI für Mikrostrukturphysik, Halle, 06120, Germany

In ultrathin films, competing exchange, anisotropy and dipolar couplings lead to a large variety of magnetic structures. These structures are obtained by means of Monte Carlo (MC) simulations at different temperatures. Three classes of magnetic structures are investigated: (i) In the Ising case, stripes, chevry structures and mazes occur. (ii) Without uniaxial anisotropy, XY-spin structures with complex arrangements of vortices are found. (iii) For intermediate uniaxial anisotropy, a spin reorientation transition occurs from in-plane to out-of-plane spins in a twisted spin phase where both mazes and vortices appear. Magnetic hysteresis cycles are also obtained with evidence for a strong directional dependence. Domain nucleation and motion are also accessible by MC relaxation at the coercive field with 360^{∘} solitons in the weak anisotropy situation and bubbles in the case of strong uniaxial anisotropy.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : oral

Corresponding author : J.-C.S. Lévy

Address for correspondence : LPTMC, case 7020, Université Paris 7, 75251 Paris cedex 05, France

Email address : levy@ccr.jussieu.fr